OpenStudy (caishaax3):

Multiply x^1/3 times x^3/7 .

OpenStudy (caishaax3):

do i need a common denominator?

OpenStudy (caishaax3):

or do i just multiply them

OpenStudy (anonymous):

\[\frac{ x^1 }{ 3 } \times \frac{ x^3 }{ 7 }\] you just multiply it.

jimthompson5910 (jim_thompson5910):

Use this rule \[\LARGE x^{a} * x^{b} = x^{a+b}\] to get \[\LARGE x^{1/3} * x^{3/7} = x^{1/3+3/7} = ??\]

jimthompson5910 (jim_thompson5910):

to add 1/3 to 3/7, yes you will need a common denominator

OpenStudy (caishaax3):

so it ends up x^3/21

OpenStudy (caishaax3):

if i just multiply it

OpenStudy (anonymous):

\[x ^{a}\times x ^{b}=x ^{a+b}\]

OpenStudy (anonymous):

a=1 b=3

OpenStudy (anonymous):

try again

OpenStudy (caishaax3):

x^1 times x^3 = x^4?

OpenStudy (anonymous):

yes

OpenStudy (caishaax3):

now what ? @Shalante

OpenStudy (anonymous):

You thought the answer was x^3/21 earlier but it is x^4/21 The denominator you multiplied was correct, but your numerator was not. Now both are correct.

OpenStudy (unklerhaukus):

\[ x^{1/3}\times x^{3/7}=x^{1/3+3/7}=x^{7/21+9/21}=x^{(7+9)/21}\]